OpenStudy (anonymous):

Show that If the limit as n approaches infinity of a(2n) =L and limit as n approaches infinity of a(2n+1)=L a(n) is convergent and the limit =L The parentheses notate subscript

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} a_{2n}=L\] \[\lim_{n \rightarrow \infty} a_{2n+1}=L\]

OpenStudy (anonymous):

Yes, that would be correct format for my question, perhaps if you could explain this definition it may help me solve the question myself.